OpenStudy (anonymous):

Water is pumped into an empty tank at a rate of r(t)=20e^(.02t) gallons per minute. Approximately how many gallons of water have been pumped into the tank in the first five minutes? Answer: 105 gal Please explain how to come to this answer and with this problem you may use a calculator.

zepdrix (zepdrix):

\[\huge\rm r(\color{orangered}{t})=20e^{.02\color{orangered}{t}}\]At time t=5 minutes we have,\[\huge\rm r(\color{orangered}{5})=20e^{(.02\cdot\color{orangered}{5})}\]Throw this into your calculator, this will complete the first half of the problem, figuring out the `rate` per minute.

zepdrix (zepdrix):

Oh the rate is variable.. hmm So are we supposed to do integration for this?

OpenStudy (whpalmer4):

I think you have to do the integral, unless you are willing to rely on the fact that the rate is pretty close to constant over that range...

OpenStudy (whpalmer4):

r(0) = 20, r(5) = 22.1034 You could use the midpoint value and be well within 1% error compared to the integral

OpenStudy (whpalmer4):

To do it with the integral, you're just integrating the rate equation times \(dt\) \[V=\int _0^5r(t) dt=\int_0^520e^{0.02t}\,dt =\]

OpenStudy (anonymous):

Hmm...I just tried to do that but it didn't come out right. I got 58.9653. Do I need to do something with this value?

OpenStudy (whpalmer4):

show me your work...what do you get for the indefinite integral?

OpenStudy (anonymous):

\[\int\limits_{0}^{5}20e^.02t\] \[20\int\limits_{0}^{5}e^.02dt\] u=.02t dt=.02dt \[20 \times.4 \int\limits_{0}^{5}e^udu\] \[.4(e^u) from 0 \to 5\] .4(e^5-e^0) = 58.9653

OpenStudy (anonymous):

for that 3rd integral that should be .02 instead of .4.

OpenStudy (whpalmer4):

Uh, no. \[u = 0.02t\]\[du = 0.02\,dt\]\[dt = \frac{du}{0.02}\] \[20\int e^{0.02t}\,dt = 20\int e^u du = \frac{20}{0.02}\int e^u du = 1000e^u = 1000e^{0.02t}\]

OpenStudy (whpalmer4):

This is a pretty common mistake, I'm afraid...but that's good, because if you remember not to make it, you eliminate a lot of mistakes :-)

OpenStudy (anonymous):

hmm I was taught a different way but i see what you did. But if i plug in my 5 it doesn't come out right so i'm kinda confused.

OpenStudy (whpalmer4):

Well, there's a right way, and a wrong way. This is a right way :-)

OpenStudy (whpalmer4):

Remember, you need to change the limits of integration if you change the variable you are integrating

OpenStudy (anonymous):

Oh right. Let me see if that works. :)

OpenStudy (whpalmer4):

But evaluating the problem with my indefinite integral: \[1000e^{0.02*5}-1000e^{0.02*0} = 1000e^{0.1}-1000e^0 \approx 1105.171 - 1000(1) = 105.171\]

OpenStudy (anonymous):

Oh my goodness I got it! :) Thank you so much

OpenStudy (whpalmer4):

Glad I could help. By the way, if you look at what @zepdrix proposed doing initially, it actually comes out quite close to the right answer! \[r(5) = 20e^{0.02*5} = 22.1034\] Using that rate for 5 minutes, we get \[V = 5*22.1034 = 110.517\] Not bad for an estimate... If we do as I suggested, and use the rate halfway through: \[r(2.5) = 20e^{0.02*2.5} = 21.0254\] \[V = 5*21.0254 = 105.127\] \[\frac{105.127-105.171}{105.171}*100\% = -0.042\%\] Close enough for government work :-)